A bullet leaves a gun with a horizontal velocity of 400 m/s. Find the horizontal and vertical displacement after 3 seconds.

  1. A bullet leaves a gun with a horizontal velocity of 400 m/s. Find the horizontal and vertical displacement after 3 seconds.

Grade:11

3 Answers

Nkurunziza
3 years ago
As the bullet moves horizontal it means the vertical displacement is zero 
v=400m/s , t= 3 s
d=vt 
=400×3
=1200m
 
 
saransh jain
3 years ago
As the bullet moves horizontal it means the vertical displacement is zero so:
Given:
v=400m/s 
t= 3 s
so
displacement = velocity * time 
=400×3
=1200m
Shivam mishra
2 years ago
Yes this is right answer of this qn.. v=400m/s 
t= 3 s
so
displacement = velocity * time 
=400×3
=1200m